AVAC is the only pharmaceutical firm producing a Vaccine.


The Demand Curve for its product is Qd = 250 – 50 P        

where P is Price and Q are packs of vaccines in ‘000


                       Total Cost Function estimated by the firm is    TC = 15 + 0.5Q     

where Q is monthly output.


a.      What is the market structure of AVAC? State its characteristics.  (1 Mark)


b.     To maximize profit,           (2 Marks)

(i)                What will be the optimum price and how many packs of Vaccine should the firm produce and sell per month?

(ii)             If this number of packs is produced and sold, what will be the firm’s monthly profit?                                                                       


c.      Using available information, draw  AVAC’s demand, marginal revenue and marginal cost curves in a graph and clearly label thefirm’s profit maximizing price, quantity and profit. Do you observe any welfare loss? If so, also indicate and label the area on the graph.(2 Marks)


d.     Assume all other pharmaceutical firms in the market start producing the Vaccine and the market becomes competitive. What will be the impact on price and marginal revenue?Would the market structure of the firm remain the same? Support your answer with help of the firm’s graph.(2 Marks)



1
Expert's answer
2021-02-01T07:55:27-0500


(A)The market structure is a monopoly.

Characteristics involve: There is a single seller in the market,the product has no close substitute and there are barriers to entry in the market.

(B)Optimum price

MR = MC,

MR = TR'(Q) = 5 - 0.04Q,

MC = TC'(Q) = 0.5.

5 - 0.04Q = 0.5,

0.04Q = 4.5,

Q = 112.5 units,

P = 5 - 0.02*112.5 = 2.75

(ii) profit

TP = 2.75*112.5 - (15 + 0.5*112.5) = 238.125

(C)






(D) Price will fall as marginal revenue decreases. The market structure won't remain the same due to increased competition.